If the magnetic field due to a long, straight current-carrying wire has a magnitude B at a distance R from the wire’s center, how far away must you be (in terms of R) for the magnetic field to decrease to B/3?
Answer #1

here,

magnetic field B due to a straight current carrying wire is B = uoi/2pIR

here   let B1 = B at R1 = R

then

B2 = B/3

so R2 = ?

so

B1R1 = B2 R2

R2 = (B * R) /(B/3)

R2 = 3 times of R is the distance

so 3R is the answer